# Off-Path TCP Hijacking in Wi-Fi Networks: A Packet-Size Side Channel Attack

### v1

@@ 4.3.3 Inferring an Acceptable Acknowledgment Number
To infer an acceptable acknowledgment number, the attacker firstly leverages the challenge ACK mechanism to locate the lower boundary of the challenge window. Then the attacker can obtain an acceptable acknowledgment number easily, i.e., add $2^{31}$ to the lower boundary.
The challenge window for TCP segment acknowledgment number is defined in RFC 5961 [56] (see Sec 2.2). As outlined in RFC 5961, the acknowledgment number space can be divided into three distinct cases, as shown in Figure 5. (i) The acknowledgment number falls within the challenge window, defined as ($SND.UNA-2G,SND.UNA-SND.WND$). (ii) The acknowledgment number resides within the acceptable ACK window, encompassing ($SND.UNA-SND.WND,SND.NXT$). (iii) Invalid acknowledgment numbers are those that exceed $SND.NXT$, denoted as $SEG.ACK>SND.NXT$.
@@
To locate the server’s challenge ACK window, the attacker impersonates the client and sends forged ACK packets to the server. The forged ACK packets carry the guessed acknowledgment numbers, as well as a sequence number in the server’s acceptable window inferred in the previous step. If the attacker sniffs a returned 68-byte encrypted frames in the Wi-Fi channel, it indicates that the guessed acknowledgment number falls within the receiver’s challenge window (as shown in step 4 of Figure 3(b)). Typically, the window size of challenge ACK is between $2^{30}$ and $2^{31}$, i.e., the challenge window is a quarter of the entire acknowledgment number range. Hence, the attacker can divide the acknowledgment range into four blocks and try at most four times to find an acknowledgment number ($ack\_challenge$) that is located in the challenge window.
@@ 5 Case Study Attacks
Similar to the prior works [14, 21], our attack focuses on compromising long-lived TCP connections. In this section, we demonstrate two cases, i.e., SSH DoS and web manipulation, to illustrate how TCP connections can be hijacked by exploiting the encrypted frame size in Wi-Fi networks. In summary, an off-path attacker can reset an SSH service within 270 seconds and inject malicious data into a HTTP web page [^4] HTTPS can prevent attackers from injecting malicious data. However, reports on HTTPS adoption [64] indicate that there are 17% of websites still based on HTTP as of August 2023. within 335 seconds.
@@ 5.1 TCP DoS Attack
Experimental Setup. This case involves three hosts: an SSH server (a rented VPS) running OpenSSH 8.4 and OpenSSL 1.1.1, a victim client (our laptop) running MacOS, and an attacker equipped with Kali 12.04 and multiple wireless network interface cards. The victim client is a supplicant in our Wi-Fi network and connects to the remote SSH server. The client will send commands to the server intermittently. Note that although the attacker and the victim supplicant are in the same Wi-Fi network, the attacker does not know the session key between the victim supplicant and the AP. The attacker attempts to terminate the connection by impersonating the victim supplicant and sending forged RST packets to the server. Taking into account the potential impact of packet transmission delays and Linux kernel versions, we strategically position servers with diverse Linux kernel versions in different locations. Detailed configuration information for these servers is provided in Table 2.
Attack Procedure. In this attack, the off-path attacker needs to infer the 4-tuple of [client IP address, client port number, server IP address, server port number] and the exact sequence number of the target TCP connection. The server’s IP address and port number are publicly known to the attacker [58, 14], thus it only needs to identify the other three remaining elements to proceed with the attack. The attacker first probe the victim client’s IP address and MAC address. Then, the attacker exploits the challenge ACK mechanism to determine the client port number of the TCP connection. Third, the attacker infers the exact sequence number by leveraging the TCP header options. Finally, a crafted RST packet carrying the inferred value is issued to the server, and the server will be tricked into terminating the current SSH connection with the victim client.
@@
Results Evaluation. Table 2 displays the outcomes of our experiments, revealing that our attack is robust to transmission delays and effective for different Linux versions. In particular, our attack exhibits low bandwidth consumption (i.e., no more than 2.3 KB/s) and maintains an average execution time of 265.9 seconds. The empirical time cost distribution is shown in Figure 6. Our attack achieves a success rate of 82% on average. For the unsuccessful attempts, the primary cause is wireless interference, leading to the attacker missing crucial encrypted frames belonging to the victim. These frames contain the server’s responses to the probe packets. We will discuss wireless interference in depth in Sec 7.1.
@@ 5.2 TCP Manipulation Attack
TCP connection hijacking poses a substantial threat to higher-layer applications, enabling malicious activities such as injecting harmful data into HTTP websites. In our research, we chose four categories, i.e., government, universities, banks, and financial institutions, which may incur significant impacts, from the Tranco [44] top one million websites based on the keywords of domain names (e.g., gov, edu, bank, and stock, etc). Subsequently, we randomly selected 1000 HTTP sites from these four categories and identified 63 vulnerable websites [^5] We consider a TCP connection that transmits page data and lasts longer than the empirical average time required for our TCP hijacking attack (335 seconds) to be a valid long-lived TCP connection. A site is considered vulnerable if it maintains no more than two such valid long-lived connections. across 10 countries, encompassing 29 government websites, 17 university websites, 7 banking websites, and 10 financial institutions websites. These websites offer various online services, including government services, customer support, financial data, and more. Attacking them can spread false information and potentially lead to significant financial losses for users. As a case in point, we demonstrate that in a typical financial website scenario, an off-path attacker can hijack the underlying TCP connection, thereby tampering with real-time financial data displayed on the victim’s web page.
@@
Experimental Setup. This attack involves three hosts: a web server, a victim client (our laptop), and an off-path attacker. We use a real financial website [^6] For ethical considerations, we anonymize this financial website in the paper. Moreover, our attack do not affect the website, since we only manipulate the web cache of the client side, i.e., our controlled laptop. as the web server. This website employs HTTP to deliver real-time financial data to the client in JSON format at 60-second intervals. Before launching the attack, the attacker can explore this website to familiarize themselves with the JSON data structure. Consequently, the attacker can discern specific data within the packet, as depicted in Figure 8(a), enabling manipulation of the victim’s web page. The server maintains a single long-lived TCP connection [^7] As recommended in RFC 2616 [38], the client typically does not maintain multiple long-lived TCP connections with the server simultaneously. with the client to transmit real-time financial data. The modern browser may open multiple concurrent TCP connections along with the long TCP connection to speed up the page loading. These concurrent TCP connections are short-lived and have minimal influence on inferring the target long-lived TCP connection. The victim client browses financial information on the web page via Wi-Fi. Both the attacker and the victim client are connected to the same Wi-Fi network. The off-path attacker attempts to detect and hijack the web connection between the victim client and the server.
@@
Attack Procedure. The web connection hijacking attack consists of five steps: (i) The attacker determines the MAC address and IP address of the victim client by leveraging ARP. (ii) By exploiting the encrypted frames, the attacker detects the client’s port number to obtain the TCP 4-tuple information. (iii) The attacker infers the exact sequence number and (iv) gets an acceptable acknowledgment number. (v) The attacker impersonates the server and injects forged TCP packets with the inferred values into the victim client. Finally, the client will accept the forged TCP packets, which subsequently update the financial information on the web page.
@@
Results Evaluation. Figure 7 illustrates the time cost and bandwidth consumption during the attack. It takes an average of 197.8 seconds to identify the client port number and 72.9 seconds to find the exact sequence number. Time cost required to locate the challenge ACK window is 3.5 seconds, while finding an acceptable acknowledgment number takes 60.6 seconds. Inferring the client port number of the connection consumes approximately 59% of the total time. This is due to the large search space of potential port numbers, which requires the attacker to sequentially probe each possible port. The average duration of the entire attack is 334.8 seconds, with an average bandwidth cost of 1.77 KB/s. In this case, the attacker needs to infer an acceptable acknowledgment number, hence the success rate of this attack is lower than the TCP DoS attack but still exceeds 70%. After obtaining all the necessary information, the attacker sends forged TCP packets to the victim client and manipulates sensitive data on the web page. Figure 8(b) shows a snapshot of the manipulated web page. The attacker alters the HSI data, modifying the values from “19,250 151 0.79% 71B” to “99,999 999 9.99%  [^9] B”.
Refer to caption (a) The attacker injects malicious data into the victim’s web connection. Refer to caption (b) The attacker manipulates the HSI data presented on the victim’s web page.
@@
6 Real-world Attacks
To assess the impact of our attack, we conduct an extensive investigation on 80 real-world Wi-Fi networks and 30 popular wireless routers. Precisely, we conduct SSH DoS attack and web hijack attack on the victim (i.e., our device) in the real-world Wi-Fi networks following the experimental setup and procedure in Sec 5. The results reveal that our attack is successful [^8] We conduct 10 iterations of the SSH DoS attack and the web hijack attack on the victim. In this context, the “successful” means that these two attacks can be successfully executed at least once in the real-world Wi-Fi network. against 69 (86.25%) of the assessed Wi-Fi networks. Besides, we analyze 30 popular wireless routers and find that all the evaluated routers cannot protect the victim from our attack.
Table 3: Experimental results in 20 real-world Wi-Fi networks.
No.
SSID
AP Vendor
IPv4/IPv6
PHY model
Encryption
W
i-Fi channel
SSH DoS
Web hijack
1 Bookstore 1 ADSLR 802.11n/ac Yes 6, 161 7/10 6/10
2 Bookstore 2 HUAWEI 802.11n/ac/ax Yes 11, 44 7/10 7/10
3 Coffee Shop 1 TP-LINK 802.11n/ac Yes 6, 60 8/10 6/10
4 Coffee Shop 2 HUAWEI 802.11n/ac Yes 1, 36 8/10 7/10
5 Coffee Shop 3 Tenda 802.11n/ac Yes 4, 153 6/10 5/10
6 Restaurant 1 D-Link 802.11n/ac Yes 5, 149 7/10 5/10
7 Restaurant 2 TP-LINK 802.11n/ac Yes 1, 153 6/10 6/10
8 Restaurant 3 iKuai 802.11n/ac Yes 1, 48 5/10 3/10
9 Office building 1 TP-LINK 802.11n/ac Yes 11, 36, 40 7/10 6/10
10 Office building 2 H3C 802.11n/ac Yes 1, 48, 153 8/10 7/10
11 Office building 3 ZTE 802.11n/ac Yes 11, 60 6/10 6/10
12 Experience Store 1 HUAWEI 802.11n/ac Yes 1, 36 7/10 6/10
13 Experience Store 2 HUAWEI 802.11n/ac Yes 11, 149 7/10 6/10
14 Experience Store 3 Tenda 802.11n/ac Yes 4,153 6/10 5/10
15 Fast Food Restaurant 1 HPE 802.11n/ac/ax Yes 6, 161, 149 6/10 4/10
16 Fast Food Restaurant 2 Ruijie 802.11n/ac Yes 1,44 6/10 6/10
17 Cinema 1 H3C 802.11n/ac No 10, 149 7/10 5/10
18 Cinema 2 HUAWEI 802.11n/ac Yes 1, 157 7/10 6/10
19 Hotel 1 HUAWEI 802.11n/ac Yes 6, 44 8/10 7/10
20 Hotel 2 Ruijie 802.11n Yes 1, 11 8/10 6/10
  • ◐ means IPv4 only and ● means both IPv4 and IPv6 are supported.
6.1 Real-world Wi-Fi Networks Evaluation
The Wi-Fi scenarios we tested cover a wide range of public settings, including coffee shops, shopping malls, restaurants, hotels, cinemas, and bookstores. The experimental results illustrate that over 86% (i.e., 69 out of 80) of the evaluated Wi-Fi networks are vulnerable to our attack. Next, we elaborate on the evaluation results.
As shown in Figure 9, out of the 80 Wi-Fi networks we assessed, 74 are found to be IPv4-only networks, while the remaining 6 have IPv6 capabilities [^9] Despite the accelerated deployment of IPv6 networks in recent years [6], the adoption of IPv6 support in Wi-Fi networks remains relatively limited. This can be attributed, in part, to the lack of IPv6 support in legacy wireless routers and the limited incentive for merchants to invest in new wireless routers..
Figure 9: Attack evaluation on 80 real-world Wi-Fi networks.
The 802.11n/ac standards are predominantly utilized (73.75%) in real-world Wi-Fi networks. This indicates that these Wi-Fi networks support two frequency bands (i.e., 2.4 GHz and 5 GHz), with their physical layer models based on the 802.11n and 802.11ac standards, respectively. There is only 17.5% (14 out of 80) of the evaluated Wi-Fi networks support 802.11ax. This is consistent with our expectations, as the 802.11ax standard was defined in 2019 and would require more time for widespread deployment. Furthermore, the simultaneous utilization of two channels in Wi-Fi networks is the most prevalent case (76.25%) due to the widespread support and default configuration of dual-channel capabilities in wireless routers. In certain scenarios, such as office buildings, Wi-Fi networks employ multiple wireless channels to enhance network performance. In our study, we identify 12 Wi-Fi networks that utilize multiple wireless channels. At first glance, the usage of multiple wireless channels might appear as a minor hurdle to our attack, as the attacker needs to perform additional channel scanning to determine the specific channel employed by the victim. Indeed, the attacker can enhance the success rate by employing a channel “eviction” strategy, which will be explained in detail in Section 7.1. Additionally, we encounter several Wi-Fi networks that operate on a single channel. When questioned, network administrators cited security considerations as the rationale behind this choice, although they did not provide any further specifics.
Out of all the evaluated networks, 69 (86.25%) of them are vulnerable to our attack. Out of the 69 vulnerable Wi-Fi networks identified, 16 of them are open networks that lack any enabled security mechanisms. This means that an attacker can potentially access and view the contents of all supplicant frames transmitted on these networks. Such a breach of supplicant privacy represents a significant concern.
We elaborate more on the experimental results of 20 Wi-Fi networks in Table 3. We take the first row of Table 3 as an example to analyze the results. In our study, the SSID “Bookstore 1” indicates a Wi-Fi network that is accessible in a bookstore. It is common practice to set the Wi-Fi SSID as the organization name, which may expose the organization’s identity. Therefore, to protect anonymity, we have anonymized the Wi-Fi SSID in this paper. This bookstore’s Wi-Fi network only supports IPv4 and its AP is produced by ADSLR. This AP provides two access channels (i.e., 6 and 161) and employs the 802.11n and 802.11ac standards. The TCP connection of the victim supplicant can be hijacked using the attack we present in Sec 4. The success rates to conduct SSH DoS and web hijacking on the victim supplicant are 70% and 60%, respectively.
Within the evaluated vulnerable Wi-Fi networks, we have observed a range of success rates for our attack, varying from 30% to 80%. The principal factor influencing this variance is the heterogeneous wireless environments (e.g., different wireless interference and channel contention) encountered in real-world Wi-Fi networks, leading to varying capabilities for attacker to capture the victim’s Wi-Fi frames. Factors such as wireless interference (e.g., from microwave ovens and Bluetooth devices) and channel contention can hinder the attacker’s ability to capture the victim’s Wi-Fi frames, resulting in the failure of attacks. For instance, as illustrated in Table 3, Coffee Shop 1, situated on a university campus, experiences less wireless interference and channel contention compared to Restaurant 3, located within a large shopping mall. Consequently, the success rate of attacks in the latter Wi-Fi network is lower due to the elevated interference and contention in that environment. It is worth noting that out of the 80 Wi-Fi networks assessed, 11 were not impacted by our attack. These Wi-Fi networks enable AP isolation and the router drops ARP requests sent by the attacker, thus thwarting our attack. We will conduct a more in-depth analysis of wireless interference and AP isolation in Section 7.
Table 4: An examination of routers from 9 AP Vendors.

### v4

@@ 4.3.3 Inferring an Acceptable Acknowledgment Number
To infer an acceptable acknowledgment number, the attacker firstly leverages the challenge ACK mechanism to locate the lower boundary of the challenge window. Then the attacker can easily obtain an acceptable acknowledgment number by adding $2^{31}$ (i.e., 2G) to the lower boundary.
The challenge window for TCP segment acknowledgment number is defined in RFC 5961 [59] (see Section 2.2). As outlined in RFC 5961, the acknowledgment number space can be divided into three distinct cases, as shown in Figure 5. (i) The acknowledgment number falls within the challenge window, defined as ($SND.UNA-2G,SND.UNA-SND.WND$). (ii) The acknowledgment number resides within the acceptable ACK window, encompassing ($SND.UNA-SND.WND,SND.NXT$). (iii) Invalid acknowledgment numbers are those that exceed $SND.NXT$, denoted as $SEG.ACK>SND.NXT$.
@@
To locate the server’s challenge ACK window, the attacker impersonates the client and sends forged ACK packets to the server. The forged ACK packets carry the guessed acknowledgment number, as well as a sequence number in the server’s acceptable window inferred in the previous step. If the attacker sniffs a returned 68-byte encrypted frames in the Wi-Fi channel, it indicates that the guessed acknowledgment number falls within the receiver’s challenge window (as shown in step 4 of Figure 3(b)). Typically, the window size of challenge ACK is between $2^{30}$ and $2^{31}$, i.e., the challenge window is a quarter of the entire acknowledgment number range. Hence, the attacker can divide the acknowledgment range into four blocks and try at most four times to find an acknowledgment number ($ack\_challenge$) that is located in the challenge window.
@@
4.4 Practical Considerations
AP Isolation. It also known as client isolation, is a security policy that can be implemented in wireless networks to separate individual devices or users from each other, enhancing network security and privacy. In the Wi-Fi network with AP isolation enabled, the AP will discard ARP requests within the WLAN, preventing the attacker from obtaining the $<MAC,IP>$ address pair of the alive supplicant. In this case, the attacker can spoof the victim’s MAC address and leverage the DHCP mechanism to obtain the victim’s $<MAC,IP>$ address pair. Specifically, the attacker first sniffs the encrypted Wi-Fi frames and identifies the MAC address of the alive supplicant. Second, the attacker spoofs the victim’s MAC address to authenticate with the AP and requests to lease a private IP address. As the DHCP server guarantees not to reallocate the leased address within the requested time and attempts to return the same network address each time the client requests an address [18], the attacker will be assigned the same private IP address that the victim is leasing. Consequently, the attacker obtains the victim’s $<MAC,IP>$ address pair (as shown in Figure 6). If the victim uses Management Frame Protection (MFP), the attacker may encounter difficulties with AP authentication when spoofing the victim’s MAC address. However, prior work has shown that implementation vulnerabilities can be abused to circumvent MFP [49, 48].
Note that our attack does not require overwriting the victim’s security context to intercept the victim’s packets [48], but rather to obtain the victim’s $<MAC,IP>$ address pair. Therefore, this approach does not necessitate the AP to support pairwise master key (PMK) caching for rapid connection. With the victim’s MAC and IP address in hand, the attacker proceeds to send forged packets to the server and detect the victim’s TCP connection, as previously described. Armed with the inferred TCP connection information, the attacker can terminate or manipulate the target TCP connection. Due to AP isolation, the attacker is unable to send packets directly to the victim supplicant. Nevertheless, the attacker can opt to send malicious packets to the AP’s external IP address, which can be obtained through ICMP ping messages, as demonstrated in previous research [69]. These packets will be forwarded to the victim supplicant through the AP.
Figure 6: Snapshot of obtaining the victim’s IP address via DHCP.
Background Traffic. The background traffic may degrade the quality of the side channel (i.e., victim’s frame size), thereby impacting the effectiveness of the attack. Specifically, if the TCP packets in the background traffic have the same size as the response (e.g., challenge ACK) from the victim server, the attacker may mistakenly identify them as the actual responses. During practical attacks, the server may send empty ACK packets (such as keep-alive ACKs) to the victim supplicant. These empty ACK will interfere with the attacker’s ability to infer the port number and acknowledgment number of the TCP connection, as they share the same size as challenge ACK packets. Fortunately, the attacker has the option to leverage SACK-ACK to complete the attack, thereby bypassing the need to contend with empty ACK packets. Specifically, (i) when inferring the port number, the attacker sends two TCP packets containing data to the server, each bearing sequence numbers $seq$ and $seq+2^{31}$ respectively. If the TCP port is accurately inferred, the attacker will encounter an 80-byte encrypted frame, as one of the two packets in question is bound to elicit the server’s SACK-ACK response. Conversely, if the inference is incorrect, the attacker will not observe the 80-byte encrypted frame. (ii) Since TCP is full duplex, the attacker can utilize SACK-ACK to infer the sequence number on the client side and consequently obtain the acknowledgment number on the server side.
Shifting Receive Window. When the victim’s TCP connection carries on ongoing traffic, the acceptable sequence and acknowledgment windows will shift during the attack. Fortunately, the attack can proceed as long as the inferred sequence number and acknowledgment number fall within the sliding window. The attacker can repeatedly infer the sequence number and acknowledgment number. Even if the receive window slides quickly enough to thwart the attacker’s inference, the attacker can opt to target the other end of the TCP connection. In this scenario, the receive window adjusts at a slower pace, as demonstrated in prior research [22, 14].
@@ 5 Case Study Attacks
In this section, we demonstrate two cases, i.e., SSH DoS and web manipulation, to illustrate how TCP connections can be hijacked by exploiting the encrypted frame size in Wi-Fi networks. In summary, an off-path attacker can reset an SSH service within 19 seconds and inject malicious data into a HTTP web page [^4] HTTPS can prevent attackers from injecting malicious data. However, reports on HTTPS adoption [67] indicate that there are 15% of websites still based on HTTP as of April 2024. within 28 seconds.
@@ 5.1 TCP DoS Attack
Experimental Setup. This case involves three hosts: an SSH server (a rented VPS) running OpenSSH 8.4 and OpenSSL 1.1.1, a victim client (our laptop) running MacOS, and an attacker equipped with Kali 12.04 and multiple wireless network interface cards. The victim client is a supplicant in our Wi-Fi network and connects to the remote SSH server. The client will send commands to the server intermittently. Note that although the attacker and the victim supplicant are in the same Wi-Fi network, the attacker does not know the session key between the victim supplicant and the AP. The attacker attempts to terminate the connection by impersonating the victim supplicant and sending forged RST packets to the server. Taking into account the potential impact of Linux kernel versions, we strategically deploy servers with a variety of Linux kernel versions. Detailed configuration information for these servers is provided in Table 2.
Attack Procedure. In this attack, the off-path attacker needs to infer the 4-tuple of [client IP address, client port number, server IP address, server port number] and the exact sequence number of the target TCP connection. The server’s IP address and port number are publicly known to the attacker [61, 14, 69], thus it only needs to identify the other three remaining elements to proceed with the attack. The attacker first probes the victim client’s IP address and MAC address. Second, the attacker exploits the TCP header options to determine the client port number and infer the exact sequence number, as outlined in Section 4. Third, a crafted RST packet carrying the inferred value is issued to the server, and the server will be tricked into terminating the current SSH connection with the victim client.
@@
Results Evaluation. Table 2 displays the outcomes of our experiments, revealing that our attack is effective for different Linux versions. In particular, our attack exhibits an average bandwidth consumption of 75.76 KB/s, while maintaining an average execution time of 18.78 seconds. The empirical time cost distribution is shown in Figure 7. Our attack achieves a success rate of 84% on average. For the unsuccessful attempts, the primary cause is wireless interference, leading to the attacker missing crucial encrypted frames belonging to the victim. These frames contain the server’s responses to the probe packets. We will discuss wireless interference in depth in Section 7.
@@ 5.2 TCP Manipulation Attack
TCP connection hijacking poses a substantial threat to higher-layer applications, enabling malicious activities such as injecting harmful data into HTTP websites. As a case in point, we demonstrate that in a typical financial website scenario, an off-path attacker can hijack the underlying TCP connection, thereby tampering with real-time financial data displayed on the victim’s web page.
@@
Experimental Setup. This attack involves three hosts: a web server, a victim client (our laptop), and an off-path attacker. We use a real financial website [^5] For ethical considerations, we anonymize this financial website in the paper. Moreover, our attack do not affect the website, since we only manipulate the web cache of the client side, i.e., our controlled laptop. as the web server. This website employs HTTP to deliver real-time Bitcoin price to the client in JSON format at 5-second intervals. Before launching the attack, the attacker can explore this website to familiarize themselves with the JSON data structure. Consequently, the attacker can discern specific data within the packet, as depicted in Figure 9(a), enabling manipulation of the victim’s web page. The victim client browses financial information on the web page via Wi-Fi. Both the attacker and the victim client are connected to the same Wi-Fi network. The off-path attacker attempts to detect and hijack the TCP connection between the victim client and the server. The server maintains a single long-lived TCP connection [^6] As recommended in RFC 2616 [39], the client typically does not maintain multiple long-lived TCP connections with the server simultaneously. with the client to transmit real-time financial data. Note that the modern browser may open multiple concurrent TCP connections along with the long TCP connection to speed up the page loading. These concurrent TCP connections are short-lived and have minimal influence on inferring the target long-lived TCP connection. Even if the server maintains multiple long-lived TCP connections with the client in some cases, the attacker can infer all the TCP connections and inject malicious data.
@@
Attack Procedure. The web connection hijacking attack consists of five steps: (i) The attacker determines the MAC address and IP address of the victim client in the WLAN. (ii) By exploiting the encrypted frames, the attacker detects the client’s port number to obtain the TCP 4-tuple information. (iii) The attacker infers the exact sequence number and (iv) gets an acceptable acknowledgment number. (v) The attacker impersonates the server and injects forged TCP packets with the inferred values into the victim client. Finally, the client will accept the forged TCP packets, which subsequently update the financial information on the web page.11
@@
(a) The attacker injects malicious data into the victim’s web connection.
(b) The attacker manipulates the Bitcoin price presented on the victim’s web page.
@@
Results Evaluation. Figure 8 illustrates the time cost and bandwidth consumption during the attack. It takes an average of 10.1 seconds to identify the client port number and 8.3 seconds to find the exact sequence number. Time cost required to find an acceptable acknowledgment number takes 9.6 seconds. The average duration of the entire attack is 28 seconds, with an average bandwidth cost of 46.3 KB/s. In this case, the attacker needs to infer an acceptable acknowledgment number, hence the success rate of this attack is lower than the TCP DoS attack but still exceeds 70%. After obtaining all the necessary information, the attacker sends forged TCP packets to the victim client and manipulates sensitive data on the web page. Figure 9(b) shows a snapshot of the manipulated web page and where attacker alters the Bitcoin price.
Table 3: Details of 30 tested wireless routers.
@@
Archer AXE300 Wi-Fi 6 WPA2/WAP3 Yes TP-LINK
Archer C80 Wi-Fi 5 WPA2/WPA3 Yes TP-LINK
Archer AX10 Wi-Fi 6 WPA2/WPA3 Yes TP-LINK
@@
6 Real-world Attacks
To assess the impact of our attack, we conduct an extensive investigation on 30 popular wireless routers and 80 real-world Wi-Fi networks. We analyze 30 popular wireless routers and find that all the evaluated routers cannot protect the victim from our attack. Besides, we conduct SSH DoS attack and web hijack attack on the victim (i.e., our device) in the real-world Wi-Fi networks following the experimental setup and procedure in Section 5. The results reveal that our attack is successful [^7] We conduct 10 iterations of the SSH DoS attack and the web hijack attack on the victim. In this context, the “successful” means that these two attacks can be successfully executed at least once in the real-world Wi-Fi network. against 75 (93.75%) of the 80 assessed Wi-Fi networks.